Bobcat Hospital purchased surgery kits on the following dates and for the following amounts

Units Price Price ($)

Total ($)

January 1 beginning balance 10 15 150

March 1 purchase

30 14 420
May 1 purchase 40 14 560
July 1 purchase 10 13 130
September 1 purchase 20 17 340
November 1 purchase 10 17 170
Total 120 $1770
Ending Inventory 25

Calculate FIFO (first in, first out)

Calculate LIFO(last in, first out)

What is the Weighted Average?
Answer #1
Average cost 14.75 =1770/120
Cost of ending inventory:
FIFO (first in, first out) 425 =(10*17)+(25-10)*17
LIFO(last in, first out) 360 =(10*15)+(25-10)*14
Weighted Average method 369 or 368.75 =25*14.75